# compat=1.1.0

### v1

@@ VI Possible UV completion
In the above discussions, we considered a low energy effective framework with explicit coupling among $\chi,\Phi_{2}$ and SM neutrinos. Due to the singlet nature of $\chi,\Phi_{2}$ under the SM gauge symmetry, such a coupling is not allowed. However, it can arise at low energy limit of a UV complete theory. The same UV completion can also explain the origin of dark sector asymmetry with possible connections to the baryon asymmetry of the Universe. There have been several proposals for asymmetric DM in the literature Nussinov:1985xr ; Kaplan:2009ag ; Davoudiasl:2012uw ; DuttaBanik:2020vfr ; Barman:2021ost ; Cui:2020dly ; Falkowski:2011xh ; Patel:2022xyv ; Biswas:2018sib ; Narendra:2018vfw ; Nagata:2016knk ; Arina:2011cu ; Arina:2012fb ; Arina:2012aj ; Narendra:2019cyt ; Mahapatra:2023dbr ; Borah:2022qln ; Borah:2023qag ; Borah:2024wos . We discuss a simple scenario where singlet right handed neutrinos are included in the SM Falkowski:2011xh . If right handed neutrino (RHN) is heavy with a bare mass term of Majorana type, then we have the canonical type-I seesaw Minkowski:1977sc ; GellMann:1980vs ; Mohapatra:1979ia ; Schechter:1980gr ; Schechter:1981cv together with asymmetric DM Falkowski:2011xh . The relevant Lagrangian for RHN denoted by $N$ and dark sector particles $\chi,\Phi_{2}$ can be written as
$-\mathcal{L}\supset y_{D}\overline{L}\tilde{H}N+y_{\chi}\overline{\chi}\Phi_{2}N+\frac{1}{2}\,M_{N}\,\overline{N^{c}}N+m_{\chi}\,\overline{\chi}\,\chi+{\rm h.c.}$ (13)
The with two or more copies of RHNs, the CP violating out-of-equilibrium decay of the lightest RHN $N\rightarrow LH,N\rightarrow\chi\Phi_{2}$ lead to production of asymmetries in lepton and dark sectors respectively. Depending upon the parameters, it is possible to have these two asymmetries differ by order of magnitudes Falkowski:2011xh . Therefore, the lepton sector asymmetry can be small which gets converted into the observed baryon asymmetry via electroweak sphalerons Kuzmin:1985mm . On the other hand, DM can have a large asymmetry till it gets partially transferred to the neutrino sector below sphaleron decoupling temperature. After the electroweak symmetry breaking, the heavy RHN mixes with the active neutrinos and which opens up a effective interaction $y_{1}\overline{\chi}\Phi_{2}\nu$, where $y_{1}\sim y_{\chi}\theta_{\nu N}$ with $\theta_{\nu N}$ being the active-sterile neutrino mixing angle. While we outline only one such UV completions, we can also consider other scenarios based on other seesaw models like type-II seesaw Mohapatra:1980yp ; Schechter:1981cv ; Wetterich:1981bx ; Lazarides:1980nt ; Brahmachari:1997cq , type-III seesaw Foot:1988aq , radiative seesaw Tao:1996vb ; Ma:2006fn or other ways of generating asymmetries like Affleck-Dine mechanism Affleck:1984fy . While these alternate UV completions have different phenomenological consequences, we confine ourselves to the low energy limit only in this work.
@@ VII Conclusion
We have proposed a mechanism to produce large neutrino asymmetry at low scale from dark matter decay. While dark matter is cosmologically stable, it can decay in the early Universe due to finite temperature effects and if DM is asymmetric like the visible matter, its decay into light neutrinos can transfer some of the dark sector asymmetries into neutrinos. The decay of DM is enabled in the vicinity of a first order phase transition which leads to sharp change in one of the dark sector particle’s mass. The large neutrino asymmetry can have interesting cosmological consequences like altering light nuclei abundance which can solve the recently reported Helium-4 anomaly. Such asymmetry can also lead to observable $\Delta N_{\rm eff}$ at future CMB experiments. In order to be in agreement with the observed baryon asymmetry, such decay of DM and hence the FOPT are required to occur below the electroweak scale. This keeps the corresponding gravitational wave peak frequencies in nHz-mHz ballpark. Due to the presence of new physics below the electroweak scale, the model can have other detection prospects which we do not discuss here. Asymmetric DM with new gauge interactions can give rise DM self-interactions motivated from small-scale structure problems of cold DM. Additionally, such gauge portal interactions also open up interesting direct detection prospects which we discuss briefly in appendix B.
@@ Appendix A Details of FOPT and Gravitational Waves
$V(\Phi_{1},\Phi_{2})$ $=\lambda_{1}\left(|\Phi_{1}|^{2}-\frac{v_{D}^{2}}{2}\right)^{2}+\mu_{\Phi_{2}}^{2}|\Phi_{2}|^{2}+\lambda_{2}|\Phi_{2}|^{4}+\lambda_{\Phi_{1}\Phi_{2}}|\Phi_{1}|^{2}|\Phi_{2}|^{2}.$ (14)
Here, $v_{D}$ is the vacuum expectation value (VEV) of the singlet scalar $\Phi_{1}(\equiv(\phi_{1}+v_{D}+i\phi_{g})/\sqrt{2})$. The Coleman-Weinberg (CW) potential Coleman:1973jx can be written as
$V_{\rm CW}(\phi_{1})=\frac{1}{(8\pi)^{2}}\sum_{i=\Phi_{1},\Phi_{2},Z_{D}}n_{i}M_{i}^{4}(\phi_{1})\left\{\log{\left(\frac{M_{i}^{2}(\phi_{1})}{v_{D}^{2}}\right)}-C_{i}\right\}$ (15)
where, $n_{\Phi_{1}}=2$, $n_{\Phi_{2}}=2$ and $n_{Z_{D}}=3$ and $C_{\Phi_{1},\Phi_{2}}=\frac{3}{2}$, $C_{Z_{D}}=\frac{5}{6}$. The physical field-dependent masses of particles coupled to $\Phi_{1}$ are
$M_{\Phi_{2}}^{2}(\phi_{1})=\mu_{\Phi_{2}}^{2}+\frac{\lambda_{\Phi_{1}\Phi_{2}}\phi_{1}^{2}}{2}\hskip 14.22636pt{\rm and}\hskip 14.22636ptM_{Z_{D}}^{2}(\phi_{1})=g_{D}^{2}\phi_{1}^{2}.$
@@
$V_{T}(\phi_{1},T)=\sum_{i=\Phi_{1},\Phi_{2},Z_{D}}\frac{n_{i}T^{4}}{2\pi^{2}}J_{B}\left(\frac{M_{i}^{2}(\phi_{1})}{T^{2}}\right)$ (16)
@@
$V_{\rm daisy}(\phi_{1},T)=-\frac{T}{2\pi^{2}}\sum_{i=\Phi_{1},\Phi_{2},Z_{D}}n_{i}[M_{i}^{3}(\phi_{1},T)-M_{i}^{3}(\phi_{1})]$ (17)
where, $M_{i}^{2}(\phi_{1},T)$=$M_{i}^{2}(\phi_{1})$ + $\Pi_{i}(T)$ and the relevant thermal masses are
$M_{\Phi_{1}}^{2}(\phi_{1},T)=M_{\Phi_{1}}^{2}(\phi_{1})+\left(\frac{\lambda_{\Phi_{1}}}{4}+\frac{\lambda_{\Phi_{1}\Phi_{2}}}{12}+\frac{g_{D}^{2}}{4}\right)T^{2},\hskip 11.38092pt$
$M_{\Phi_{2}}^{2}(\phi_{1},T)=M_{\Phi_{2}}^{2}(\phi_{1})+\left(\frac{\lambda_{\Phi_{2}}}{4}+\frac{\lambda_{\Phi_{1}\Phi_{2}}}{12}+\frac{g_{D}^{2}}{4}\right)T^{2},$
$M_{Z_{D}}^{2}(\phi_{1},T)=M_{Z_{D}}^{2}(\phi_{1})+g_{D}^{2}T^{2}.$
@@
The above effective potential shows different profiles depending on the temperature. When the temperature reaches a critical point, the potential develops two identical minima separated by a barrier. Below the critical temperature, the non-zero minimum becomes the true vacuum, while the zero minimum becomes the false vacuum. Due to the presence of the barrier, the Universe remains in the false vacuum, and then transitions from the false vacuum to the true vacuum by tunneling through the barrier. The rate of tunneling per unit time per unit volume can be estimated as $\Gamma(T)=\mathcal{A}(T)e^{-S_{3}(T)/T},$ where $\mathcal{A}(T)\sim T^{4}$ and $S_{3}$ is Euclidean action. Here, the Hubble parameter is given by ${\bf\mathcal{H}}(T)\simeq 1.66\sqrt{g_{*}}T^{2}/M_{\rm Pl}$ with $g_{*}$ being the dof of the radiation component. The energy difference between the true and the false vacuum is $\Delta V_{\rm eff}\equiv V_{\rm eff}(\phi_{\rm false},T)-V_{\rm eff}(\phi_{\rm true},T).$ The amount of vacuum energy released during FOPT can parameterized as $\alpha_{*}=\frac{\epsilon_{*}}{\rho_{\rm rad}}$, where radiation energy density of the Universe, $\rho_{\rm rad}=g_{*}\pi^{2}T^{4}/30$ and $\epsilon_{*}=\left[\Delta V_{\rm eff}-\frac{T}{4}\frac{\partial\Delta V_{\rm eff}}{\partial T}\right]_{T=T_{n}}.$ The inverse duration of FOPT is defined as $\frac{\beta}{{\mathcal{H}}(T)_{*}}\simeq T\frac{d}{dT}\left(\frac{S_{3}}{T}\right)|_{T=T_{n}}.$ In this work to calculate the Euclidean action, we fitted the effective potential to a generic potential for which action calculations are described in Adams:1993zs . This prescription of action estimation is described with details in Borah:2022cdx .
@@
$v_{w}=\begin{cases}\sqrt{\frac{\Delta V_{\rm eff}}{\alpha_{*}\rho_{\rm rad}}}&\text{if}\,\,\sqrt{\frac{\Delta V_{\rm eff}}{\alpha_{*}\rho_{\rm rad}}}<v_{J}\\ 1&\text{if}\,\,\sqrt{\frac{\Delta V_{\rm eff}}{\alpha_{*}\rho_{\rm rad}}}\geq v_{J}.\\ \end{cases}$ (23)

### v2

@@ VI Possible UV completion
In the above discussions, we considered a low energy effective framework with explicit coupling among $\chi,\Phi_{2}$ and SM neutrinos. Due to the singlet nature of $\chi,\Phi_{2}$ under the SM gauge symmetry, such a coupling is not allowed. However, it can arise at low energy limit of a UV complete theory. The same UV completion can also explain the origin of dark sector asymmetry with possible connections to the baryon asymmetry of the Universe. There have been several proposals for asymmetric DM in the literature Nussinov:1985xr ; Kaplan:2009ag ; Davoudiasl:2012uw ; DuttaBanik:2020vfr ; Barman:2021ost ; Cui:2020dly ; Falkowski:2011xh ; Patel:2022xyv ; Biswas:2018sib ; Narendra:2018vfw ; Nagata:2016knk ; Arina:2011cu ; Arina:2012fb ; Arina:2012aj ; Narendra:2019cyt ; Mahapatra:2023dbr ; Borah:2022qln ; Borah:2023qag ; Borah:2024wos .
For simplicity, here we consider a low scale seesaw scenario based on inverse seesaw Mohapatra:1986aw ; Mohapatra:1986bd ; Gonzalez-Garcia:1988okv where the desired dark fermion coupling to active neutrinos via active-sterile neutrino mixing can be realised while being consistent with light neutrino mass and sub-dominant washout of dark fermion asymmetry. Inverse seesaw scenario requires two types of sterile fermions $N,S$ with the following Lagrangian
$\mathcal{L}\supset y_{D}\overline{L}\tilde{H}N+M_{S}\overline{N}S+\frac{1}{2}\,\mu_{S}\,\overline{S^{c}}S+{\rm h.c.}$ (13)
Here $\mu_{S}$ is the only lepton number violating term which can be naturally small. After electroweak symmetry breaking, we have a Dirac mass term $M_{D}=\frac{y_{D}v}{\sqrt{2}}$ with $v$ being the VEV of the SM Higgs doublet $H$. In the limit $\mu_{S}\ll M_{D}\ll M_{S}$, light neutrino mass at leading order is given by
$m_{\nu}=M_{D}M^{-1}_{S}\mu_{S}(M^{T}_{S})^{-1}M^{T}_{D}.$ (14)
Thus, smallness of $\mu_{S}$ keeps active neutrinos light while also allowing large active-sterile mixing
$\theta^{2}_{\nu N}=\frac{m_{\nu}}{\mu_{S}}.$ (15)
In order to generate effective $\chi-\nu$ coupling at low energy (given by Eq. (2)), we couple the RHN $N$ and dark sector particles $\chi,\Phi_{2}$ as
$\mathcal{L}\supset y_{\chi}\overline{\chi}\Phi_{2}N+{\rm h.c.}$ (16)
After the electroweak symmetry breaking, the heavy RHN mixes with the active neutrinos which opens up an effective interaction $y_{1}\overline{\chi}\Phi_{2}\nu$, where $y_{1}\sim y_{\chi}\theta_{\nu N}$ with $\theta_{\nu N}$ being the active-sterile neutrino mixing angle given by Eq. (15). Given an alternate source of dark sector asymmetry, DM can carry this large asymmetry until it gets partially transferred to the neutrino sector below sphaleron decoupling temperature.
In order to generate the large initial dark sector asymmetry, we outline a setup based on the Affleck-Dine (AD) mechanism Affleck:1984fy . While the same mechanism can also explain the observed baryon asymmetry of the Universe, we focus only on the aspect of generating large neutrino asymmetry, which is the focus of this work. Consider an Affleck-Dine field $\Phi$ singlet under the SM symmetries but charged under a dark global $U(1)_{D}$ symmetry. Let the $U(1)_{D}$ charge of $\Phi$ be -2 while that of dark matter $\chi$ is 1. The relevant part of the Lagrangian involving $\Phi$ is
$\mathcal{L}\supset-(Y_{D}\overline{\chi^{c}}\chi\Phi+{\rm h.c.})-V(\Phi),$ (17)
where
$V(\Phi)=m_{\Phi}^{2}|\Phi|^{2}+\lambda_{\Phi}|\Phi|^{4}+(\epsilon m_{\Phi}^{2}\Phi^{2}+{\rm h.c.})$ (18)
For $\epsilon\neq 0$, the last term in the potential $V(\Phi)$ breaks global $U(1)_{D}$ explicitly. While there can be other terms too breaking $U(1)_{D}$ explicitly, the quadratic term used here helps to obtain an analytical expression for the asymmetry Mohapatra:2021aig . For small explicit $U(1)_{D}$ breaking term, we have an approximately conserved Noether’s current given by
$J_{\mu}=iQ_{\Phi}[\Phi^{\dagger}\partial_{\mu}\Phi-(\partial_{\mu}\Phi^{\dagger})\Phi],$ (19)
with $Q_{\Phi}$ being the global charge of $\Phi$. Using $\Phi=(\phi_{1}+i\phi_{2})/\sqrt{2}$, the number density of the approximately conserved charge is
$n_{\Delta\chi}=Q_{\Phi}(\dot{\phi_{1}}\phi_{2}-\phi_{1}\dot{\phi_{2}}).$ (20)
Due to this explicit violation, a non-zero lepton asymmetry is created through the cosmic evolution of $\Phi$ which gets transferred to the dark sector as $\Phi$ decays into $\chi$. For $\Phi\gtrsim\frac{m_{\Phi}}{\sqrt{\lambda_{\Phi}}}(\equiv\Phi^{*})$ (where $M_{P}$ denotes reduced Planck mass), the quartic term $\lambda_{\Phi}\lvert\Phi\rvert^{4}$ dominates and $\Phi\propto 1/a$. Once $\Phi$ reaches $\Phi^{*}$, difference in the real and imaginary values of $\Phi$ creates an asymmetry in the $\Phi$ condensate which oscillates with period $T_{\rm asy}=\frac{\pi}{\epsilon m_{\Phi}}$. Now, the comoving asymmetry ($N_{\Delta\chi}(t)\equiv\left(\frac{a(t)}{a_{I}}\right)^{3}n_{\Delta\chi}(t)$, with $a_{I}$ being the scale factor at the end of inflation) generated for $t>t_{*}$ can be written as Mohapatra:2021aig ; Borah:2022qln
$N_{\Delta\chi}(t)$ $\simeq 4Q_{\Phi}\,A\,\phi_{1,I}\,\phi_{2,I}\left(\frac{\phi_{I}}{\Phi^{*}}\right)\int_{t_{*}}^{t}dt^{\prime}\,{\cos}(m_{1}(t^{\prime}-t_{*}))\,{\cos}(m_{2}(t^{\prime}-t_{*}))\,e^{-\Gamma_{\Phi}(t^{\prime}-t_{*})}$ (21)
where $\Gamma_{\Phi}$ indicates the total decay rate of the inflaton $\Phi$ to $\chi$. $\phi_{1,I}\,\phi_{2,I}$ indicate the initial values of the real and imaginary parts of $\Phi$ and $\phi_{I}=\sqrt{(\phi_{1,I})^{2}+(\phi_{2,I})^{2}}$, while $m^{2}_{1}=m^{2}_{\Phi}-2A$, $m^{2}_{2}=m^{2}_{\Phi}+2A$ with $A=\epsilon m^{2}_{\Phi}$. Asymmetry thereby created is transferred to the dark sector through decay $\Phi\rightarrow\chi\chi$. Under the assumption of $2A\gg\Gamma_{\Phi}m_{\Phi}$, for $t\gtrsim 1/\Gamma_{\Phi}$, the integral above simplifies to [^5] Here, we have used the analytical expression derived in Mohapatra:2021aig $\mathcal{I}\equiv\int_{t_{*}}^{t}dt^{\prime}\,{\cos}(m_{1}(t^{\prime}-t_{*}))\,{\cos}(m_{2}(t^{\prime}-t_{*}))\,e^{-\Gamma_{\Phi}(t^{\prime}-t_{*})}\simeq\frac{\gamma}{2m_{\Phi}}\left(\frac{1}{2+\gamma^{2}-2\sqrt{1-4\epsilon^{2}}}+\frac{1}{2+\gamma^{2}+2\sqrt{1-4\epsilon^{2}}}\right)$ and assume $2\epsilon\gg\gamma$ or equivalently, $2A\gg\Gamma_{\Phi}m_{\Phi}$ to maximize the asymmetry.
$N_{\Delta\chi}(t)$ $\simeq C\frac{\gamma}{8\epsilon^{2}m_{\Phi}},$ (22)
where $\gamma=\Gamma_{\Phi}/m_{\Phi}$ and $C=4Q_{\Phi}\,A\,\phi_{1,I}\,\phi_{2,I}\left(\frac{\phi_{I}}{\Phi^{*}}\right)$. Also, we have $\gamma\ll 1$ for narrow width and $\epsilon\ll 1$ for soft breaking of global symmetry. In Fig. 6, we show the evolution of the comoving asymmetry $N_{\Delta\chi}(t)$. The asymmetry initially rises from zero and then oscillates until $t\gtrsim 1/\Gamma_{\Phi}$, when its amplitude exponentially damps to reach the constant value given by Eq. (22). The final constant value of the asymmetry is consistent with the neutrino asymmetry required to solve the recently reported Helium-4 anomaly.
Figure 6: Evolution of dark sector asymmetry with time for a benchmark choice of model parameters. The final asymmetry is of the same order as the one required to address the Helium-4 anomaly.
If the same decay of the AD field $\Phi$ to dark matter also reheats the Universe to a temperature $T_{\rm rh}$, then we can relate its decay width $\Gamma_{\Phi}$ to the reheat temperature $T_{\rm rh}\simeq\sqrt{\Gamma_{\Phi}M_{\rm P}}$. The lepton asymmetry initially rises from zero and then oscillates until $t\gtrsim 1/\Gamma_{\Phi}$, when its amplitude exponentially damps to reach the constant value given by Lloyd-Stubbs:2020sed ; Mohapatra:2021aig ; Borah:2022qln
$Y^{\rm in}_{\Delta\chi}=\frac{(n_{\chi}-n_{\bar{\chi}})}{s}\Big|_{T_{\rm rh}}=\frac{N_{\Delta\chi}(T_{\rm rh})}{s(T_{\rm rh})}\left(\frac{a_{\rm I}}{a_{\rm rh}}\right)^{3}=\frac{N_{\Delta\chi}(T_{\rm rh})}{s(T_{\rm rh})}\left(\frac{a_{\rm I}}{a_{*}}\right)^{3}\left(\frac{a_{*}}{a_{\rm rh}}\right)^{3}.$ (23)
Using $a\propto 1/\Phi$ for $a<a_{*}{}$ and $a\propto t^{2/3}\propto\mathcal{H}^{-2/3}$ for $a_{*}<a<a_{\rm rh}$, we obtain
$Y^{\rm in}_{\Delta\chi}\simeq\frac{N_{\Delta\chi}(T_{\rm rh})}{s(T_{\rm rh})}\left(\frac{\Phi_{*}}{\phi_{I}}\right)^{3}\left(\frac{\mathcal{H}_{\rm rh}}{\mathcal{H}_{*}}\right)^{2}.$ (24)
Using $\mathcal{H}^{2}_{\rm rh}=\frac{\pi^{2}}{90}g_{*}\frac{T^{4}_{\rm rh}}{M^{2}_{P}}$, $\mathcal{H}^{2}_{*}=\frac{m^{2}_{\Phi}\Phi_{*}^{2}}{6M^{2}_{P}}$ and the expression for $N_{\Delta\chi}(T_{\rm rh})$ given in Eq. (22), we obtain
$Y^{\rm in}_{\Delta\chi}\simeq\frac{3}{4}\sqrt{\frac{\pi^{2}}{90}g_{*}}Q_{\Phi}\phi_{1,I}\phi_{2,I}\left(\frac{1}{\phi_{I}}\right)^{2}\frac{T_{\rm rh}^{3}}{\epsilon m_{\Phi}^{2}M_{\rm P}}\simeq\frac{3}{8}\sqrt{\frac{\pi^{2}}{90}g_{*}}Q_{\Phi}\frac{T_{\rm rh}^{3}}{\epsilon m_{\Phi}^{2}M_{\rm P}}\sin(2\theta_{\phi}).$ (25)
Here $\theta_{\phi}$ is the phase of $\Phi$, which we can consider to be of $\mathcal{O}(1)$. Now, the presence of lepton number violating interaction given by $\epsilon$ can lead to the washout of the generated asymmetry. This can happen through scatterings violating $U(1)_{D}$ charge by 4 units: $\chi\chi\leftrightarrow\overline{\chi}~\overline{\chi}$, mediated by $\Phi$ exchange and the $\epsilon$ term. If the decoupling temperature of such process is higher than the reheat temperature $T_{\rm rh}$, the washout effect would be absent. This leads to the following condition
$\Gamma_{\rm wo}(\Delta L=4)\equiv T_{\rm rh}^{3}\frac{Y_{D}^{4}\epsilon^{2}T_{\rm rh}^{2}}{64\pi m_{\Phi}^{4}}\lesssim\mathcal{H}(T_{\rm rh})\equiv\sqrt{\frac{\pi^{2}}{90}g_{*}}\frac{T_{\rm rh}^{2}}{M_{\rm P}}.$ (26)
Additionally, demanding the AD field to be not part of the thermal bath requires $T_{\rm rh}<m_{\Phi}$ which ensures washouts due to inverse decays being sub-dominant. Fig. 7 shows comoving dark asymmetry as a function of $\epsilon$ for a benchmark choice of parameters. The pink shaded region corresponds to the region where analytical approximation of asymmetry given by Eq. (22) does not apply. The entire parameter space shown in Fig. 7 correspond to sub-dominant washouts mediated by $\Phi$ or washouts due to inverse decay into $\Phi$.
Another washout of dark asymmetry can arise due to the coupling of $\chi$ with Majorana heavy fermions $N_{i}$. One can have $\Delta L=2$ washout process like $\chi\Phi^{\dagger}_{2}\leftrightarrow\overline{\chi}~\Phi_{2}$ mediated by $N_{i}$. Demanding this process to be inefficient at $T=T_{\rm rh}$ leads to
$\Gamma_{\rm wo}(\Delta L=2)\equiv T_{\rm rh}^{3}\frac{y_{\chi}^{4}\mu^{2}_{S}}{64\pi M_{S}^{4}}=T_{\rm rh}^{3}\frac{y_{1}^{4}\mu^{2}_{S}}{64\pi(\theta_{\nu N}M_{S})^{4}}\lesssim\mathcal{H}(T_{\rm rh})\equiv\sqrt{\frac{\pi^{2}}{90}g_{*}}\frac{T_{\rm rh}^{2}}{M_{\rm P}},$ (27)
where we use $y_{1}=y_{\chi}\theta_{\nu N}$. Using active-sterile mixing given by Eq. (15), this leads to the following upper bound
$T_{\rm rh}<\sqrt{\frac{\pi^{2}}{90}g_{*}}\frac{64\pi}{y^{4}_{1}}\frac{m^{2}_{\nu}}{M_{\rm P}}\left(\frac{M_{S}}{\mu_{S}}\right)^{4}.$ (28)
For $y_{1}\sim 10^{-9}$ and $m_{\nu}<0.1$ eV, $T_{\rm rh}$ can be as large as $10^{16}$ GeV for $M_{S}/\mu_{S}\sim 10^{4}$. For larger $M_{S}/\mu_{S}$ which is generic in inverse seesaw scenarios, this bound can be satisfied trivially.
In order not to overproduce baryon asymmetry of Universe, it is also required to ensure that the large dark fermion asymmetry does not get transferred to SM leptons via scatterings like $\chi\Phi^{\dagger}_{2}\rightarrow lH$ mediated by $N$. Keeping this $\Delta L=0$ conversion process out-of-equilibrium at $T=T_{\rm rh}$ leads to
$\Gamma_{\rm transfer}\equiv T_{\rm rh}^{3}\frac{y^{2}_{D}y_{\chi}^{2}}{64\pi M_{S}^{2}}=T_{\rm rh}^{3}\frac{y^{2}_{1}}{32\pi v^{2}}\lesssim\mathcal{H}(T_{\rm rh})\equiv\sqrt{\frac{\pi^{2}}{90}g_{*}}\frac{T_{\rm rh}^{2}}{M_{\rm P}},$ (29)
where we use $y_{\chi}=\frac{y_{1}}{\theta_{\nu N}}=\frac{\sqrt{2}\,y_{1}\,M_{S}}{y_{D}\,v}$.
Figure 7: Comoving dark sector asymmetry as a function of $\epsilon$ for one benchmark choice of model parameters. The analytical expression for asymmetry is not valid in the pink shaded region requiring numerical estimates.
To verify the out-of-equilibrium conditions given by Eqs. (26), (27) and (29) for the benchmark plots shown in Fig. 6 and 7, we plot the Hubble expansion rate and interaction rates for washout and transfer processes as a function of reheating temperature in Fig. 8. The region below the black solid line denotes the out-of equilibrium regime. The blue, magenta and red colored dashed lines denote the interaction rate given in Eqs. (26), (27) and (29) respectively. The rest of the parameters are kept fixed at $Y_{D}=4.2\times 10^{-6},m_{\Phi}=10^{6}$ GeV, $m_{\chi}=14$ MeV and $\epsilon=2\times 10^{-12}$ (same values as in Fig. 6 and Fig. 7). In addition, we use $y_{1}=10^{-9}$ which is consistent with left plot of Fig. 4. For the condition in Eq. (27), we consider the ratio $\frac{M_{S}}{\mu_{S}}$ to be $10^{4}$. With these, the maximum reheating temperature to keep all the washout and transfer processes out-of-equilibrium is found to be $\sim 10^{6}$ GeV.
Figure 8: Comparison of asymmetry washout rates $\Gamma_{\rm wo}(\Delta\rm L=4)$, $\Gamma_{\rm wo}(\Delta\rm L=2)$ and asymmetry transfer rate $\Gamma_{\rm transfer}$ with Hubble expansion rate $\mathcal{H}$ as a function of reheating temperature. The blue, magenta and red color dashed lines denote the interaction rate given in Eq. (26), (27) and (29) whereas black solid line denotes the corresponding Hubble expansion rate.
@@ VII Conclusion
We have proposed a mechanism to produce large neutrino asymmetry at low scale from dark matter decay. While dark matter is cosmologically stable, it can decay in the early Universe due to finite temperature effects and if DM is asymmetric like the visible matter, its decay into light neutrinos can transfer some of the dark sector asymmetries into neutrinos. The decay of DM is enabled in the vicinity of a first order phase transition which leads to sharp change in one of the dark sector particle’s mass. The large neutrino asymmetry can have interesting cosmological consequences like altering light nuclei abundance which can solve the recently reported Helium-4 anomaly. Such asymmetry can also lead to observable $\Delta N_{\rm eff}$ at future CMB experiments. In order to be in agreement with the observed baryon asymmetry, such decay of DM and hence the FOPT is required to occur below the electroweak scale. This keeps the corresponding gravitational wave peak frequencies in nHz-mHz ballpark. Due to the presence of new physics below the electroweak scale, the model can have other detection prospects the details of which we do not discuss here. We briefly comment upon interesting direct detection prospects of DM via Higgs portal interactions in Appendix B. While we have not explored the possibility of explaining the observed baryon asymmetry within the same setup, it can be generated independently from heavy right handed neutrino decay Gu:2010xc . Given the upper bound on reheating temperature $T_{\rm rh}<10^{6}$ GeV from the requirement of keeping washout and asymmetry transferring scattering processes out of equilibrium, the leptogenesis is likely to occur via resonant enhancement or non-thermally. We leave such possible extensions of the framework to future studies.
@@ Appendix A Details of FOPT and Gravitational Waves
$V(\Phi_{1},\Phi_{2})$ $=\frac{\lambda_{1}}{4}\left(\Phi_{1}^{2}-\frac{v_{D}^{2}}{2}\right)^{2}+\mu_{\Phi_{2}}^{2}|\Phi_{2}|^{2}+\lambda_{2}|\Phi_{2}|^{4}+\frac{\lambda_{\Phi_{1}\Phi_{2}}}{2}\Phi_{1}^{2}|\Phi_{2}|^{2}.$ (30)
Here, $v_{D}$ is the vacuum expectation value (VEV) of the singlet scalar $\Phi_{1}(\equiv\phi_{1})$. The Coleman-Weinberg (CW) potential Coleman:1973jx can be written as
$V_{\rm CW}(\phi_{1})=\frac{1}{(8\pi)^{2}}\sum_{i=\phi_{1},\Phi_{2},\chi}n_{i}M_{i}^{4}(\phi_{1})\left\{\log{\left(\frac{M_{i}^{2}(\phi_{1})}{v_{D}^{2}}\right)}-C_{i}\right\}$ (31)
where, $n_{\phi_{1}}=1$, $n_{\Phi_{2}}=2$, $n_{\chi}=4$ and $C_{\phi_{1},\Phi_{2},\chi}=\frac{3}{2}$. The physical field-dependent masses of particles coupled to $\Phi_{1}$ are
$M_{\phi_{2}}^{2}(\phi_{1})=\mu_{\Phi_{2}}^{2}+\frac{\lambda_{\Phi_{1}\Phi_{2}}}{2}\phi_{1}^{2},\hskip 2.84544ptM_{\chi}^{2}(\phi_{1})=(m_{0}+y_{\chi\Phi_{1}}\phi_{1})^{2}.$
@@
$V_{T}(\phi_{1},T)=\sum_{i=\phi_{1},\Phi_{2},\chi}\frac{n_{i}T^{4}}{2\pi^{2}}J_{B}\left(\frac{M_{i}^{2}(\phi_{1})}{T^{2}}\right)$ (32)
@@
$V_{\rm daisy}(\phi_{1},T)=-\frac{T}{2\pi^{2}}\sum_{i=\phi_{1},\Phi_{2}}n_{i}[M_{i}^{3}(\phi_{1},T)-M_{i}^{3}(\phi_{1})]$ (33)
where, $M_{i}^{2}(\phi_{1},T)$$M_{i}^{2}(\phi_{1})$ + $\Pi_{i}(T)$ and the relevant thermal masses are
$M_{\phi_{1}}^{2}(\phi_{1},T)=M_{\phi_{1}}^{2}(\phi_{1})+\left(\frac{\lambda_{1}}{4}+\frac{\lambda_{\Phi_{1}\Phi_{2}}}{6}\right)T^{2},\hskip 11.38092pt$
$M_{\Phi_{2}}^{2}(\phi_{1},T)=M_{\Phi_{2}}^{2}(\phi_{1})+\left(\frac{\lambda_{2}}{3}+\frac{\lambda_{\Phi_{1}\Phi_{2}}}{12}\right)T^{2}.$
@@
The above effective potential shows different profiles depending on the temperature. When the temperature reaches a critical point, the potential develops two identical minima separated by a barrier. Below the critical temperature, the non-zero minimum becomes the true vacuum, while the zero minimum becomes the false vacuum. Due to the presence of the barrier, the Universe remains in the false vacuum, and then transitions from the false vacuum to the true vacuum by tunneling through the barrier. The rate of tunneling per unit time per unit volume can be estimated as $\Gamma(T)=\mathcal{A}(T)e^{-S_{3}(T)/T},$ where $\mathcal{A}(T)\sim T^{4}$ and $S_{3}$ is Euclidean action. Here, the Hubble parameter is given by ${\bf\mathcal{H}}(T)\simeq 1.66\sqrt{g_{*}}T^{2}/M_{\rm P}$ with $g_{*}$ being the dof of the radiation component. The energy difference between the true and the false vacuum is $\Delta V_{\rm eff}\equiv V_{\rm eff}(\phi_{\rm false},T)-V_{\rm eff}(\phi_{\rm true},T).$ The amount of vacuum energy released during FOPT can parameterized as $\alpha_{*}=\frac{\epsilon_{*}}{\rho_{\rm rad}}$, where radiation energy density of the Universe, $\rho_{\rm rad}=g_{*}\pi^{2}T^{4}/30$ and $\epsilon_{*}=\left[\Delta V_{\rm eff}-\frac{T}{4}\frac{\partial\Delta V_{\rm eff}}{\partial T}\right]_{T=T_{n}}.$ The inverse duration of FOPT is defined as $\frac{\beta}{{\mathcal{H}}(T)_{*}}\simeq T\frac{d}{dT}\left(\frac{S_{3}}{T}\right)|_{T=T_{n}}.$ In this work to calculate the Euclidean action, we fitted the effective potential to a generic potential for which action calculations are described in Adams:1993zs . This prescription of action estimation is described with details in Borah:2022cdx .
@@
[…]